OpenStudy (anonymous):

The density of oxygen gas at 293K and 3 atm pressure is?

OpenStudy (unklerhaukus):

pV=nRT n/V=p/RT

OpenStudy (anonymous):

@UnkleRhaukus i got n/v = 0.124

OpenStudy (anonymous):

then wat to do?

OpenStudy (unklerhaukus):

what at the units for n/

OpenStudy (anonymous):

n = given mass/molar mass

OpenStudy (unklerhaukus):

n is the number of moles

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so n = given mass/molar mass

OpenStudy (unklerhaukus):

what is the molar mass of oxygen \(O_2\)

OpenStudy (anonymous):

32

OpenStudy (unklerhaukus):

32 grams / mol

OpenStudy (anonymous):

yup i got it it is 4g/l

OpenStudy (unklerhaukus):

to get density we want \[\rho=m/V\]mass per unit volume

OpenStudy (anonymous):

thxxx for helping!

OpenStudy (unklerhaukus):

thats ok.